# Discrete Chi-Square Method can model and forecast complex time series

### v5

@@ 3.11.3 Stability ( $\mathrm{C}_{3}$)
The artificial bootstrap data sets (Equation 31) represent “small changes in the input data”, while the bootstrap results for the model parameters represent “small chances in the solution” (Section 1: $\mathrm{C}_{3}$ condition formulation). We routinely check the stability of these solutions (e.g. Jetsu 2025, Figure S5). The unstable models, where the model parameter changes are large, are rejected (Section 2.1: $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}” models).
@@ 3.12 Forecast
There are numerous techniques for forecasting a time series [e.g., 22, 26, 31]. The DCM model $g(t_{i},\bm{\beta})$ can be used to forecast. We divide all data into the forecasting data and the forecasted data. The time points, the observations and the errors of these samples are
@@
This “Forecast-test” technique revealed at least five real signals in the sunspot record (Jetsu 2025).
@@
during the chosen sample window $\Delta T^{\prime}=t_{n}^{\prime}-t_{1}^{\prime}$. Jetsu (2025) used this $m_{\mathrm{Fore}}$ parameter to postdict the known $\Delta T^{\prime}$ time intervals of prolonged solar activity minima, like the Maunder minimum between the years 1640 and 1720. Since the known mean level of all sunspot data was $m$, Jetsu (2025) used these three criteria for correct postdictions of past prolonged activity minima $\Delta T^{\prime}$ time intervals:
@@
The analyses of all seven complex time series indicate that the relative accuracy of amplitude and period estimates are lower than the relative accuracy of signal minimum and maximum epoch estimates (Tables 1 - 7). This statistical effect is the same when DCM is applied to the time series of any arbitrary phenomenon. This effect would, for example, explain why our solar cycle amplitude forecasts are less accurate than our solar cycle minimum and maximum epoch forecasts (Jetsu 2025). DCM detects the correct simulated period values, but DFT detects less accurate period values, which are not always correct. The leakage of DFT spectral power [30, 19] would explain why correct signal periods have not been detected earlier from the sunspot data.
@@ 4 Discussion
DCM proceeds through two stages. It computes the $R$ (Equation 14) or $\chi^{2}$ (Equation 15) values for a massive number of LS fits (Equation 50). The Gauß-Markov theorem ensures that the DCM model having the lowest $R$ or $\chi^{2}$ is inevitably always found, and the result for the non-linear iteration is unique (Equation 22). This first stage devours CPU. The second fast stage, the Fisher-test, then reveals the best DCM model alternative (Equation 48). Practical time series analysis applications, like digital signal processing, require fast computational algorithms [30]. Long computation time is the main AL constraining DCM. For example, Jetsu (2025) computed about 4.5 million LS fits to search for four signals in the monthly sunspot data $(n=3287,K_{1}=4,K_{2}=1,K_{3}=0,n_{\mathrm{L}}=100,n_{\mathrm{S}}=30,n_{\mathrm{B}}=20)$. This required several months of CPU. The parallel Python code computations took a few days. Nevertheless, the CPU used is beside the point if a well-posed computational solution can be found to any challenging scientific ill-posed problem.
@@
5 Appendix
In this appendix, we show that our DCM can model and forecast the El Niño phenomenon between the years 1870 and 2024. This result may be the “holy grail” of climatology. The solar forcing most probably causes the detected $5.850\pm 0.085$, $12.82\pm 0.40$ and $19.30\pm 0.83$ year signals. These strict periodicities indicate that the solar cycle is caused by the planets. Our results confirm that the DCM is an ideal time series analysis method for detecting an unknown number of signals superimposed on an unkown trend.
The El Niño phenomenon is considered unpredictable [52, 39, 6, 25, 32, 35, 50, 40]. Our analysed data are the yearly means in the The Seasonal Mean Nino 4 HadISST1.1 (NOAA PSL) sample. [^5] The data from https://psl.noaa.gov/data/timeseries/month/data/nino4.long.anom.data were retrieved on January 29th, 2026. There were no values for the last four months of 2025. The full data contains $n=155$ yearly means between 1870 and 2024.
5.1 Trend
@@
We detect significant signals only from the weighted data. Therefore, the trend and the forecasts are solved from the weighted data. The weighted DCM analysis results for all data are given in Table S1. For one, two, three and four signals, the best trend alternative is the linear trend $K_{1}=1$. We use this $K_{1}=1$ trend in our DCM analysis of El Niño data. For the $\mathcal{M}$=3 model, the positive linear trend is $2M_{1}=0.56\pm 0.20$ $\mathrm{C}^{o}$during $\Delta T=154$ years (Equation 5), like in all $\mathcal{M}$=1-4 models.
5.2 Forecasting data
@@
We use the first half of all data $(n=78)$ to forecast the second half $(n=77)$. The results of weighted DCM analysis are given in Table S2. The periods of the three strongest signals are $5.850\pm 0.085$, $12.82\pm 0.40$ and $19.30\pm 0.83$ years (Model $\mathcal{M}$=3). The forecasts of the two and three signals models, open circles, are shown in Figures S1 and S2. For the second half of the data, the forecast $\chi^{2}$ values for the one, two, three and four signals models are 421, 336, 370 and 428, respectively. Hence, the two signal $\mathcal{M}$=2 model gives the best forecast.
5.3 All data
@@
The DCM analysis results for all weighted data are given in Table S3. The three signals $\mathcal{M}$=3 model is the best. The four signal model is unstable $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}})$. The three and four signal models are shown in Figures S3 and S4. For both alternatives, the predicted values, open circles, after 2024 are given in Table S4. Both alternatives give essentially the same forecast.
@@
For the sake of consistency, we also give the results of non-weighted analysis of all data (Table S5). We detect the same signals as from non-weighted data, but the signal significances $Q_{\mathrm{F}}$ are lower. This confirms that the data error information is crucial.
5.4 Main result
Our DCM can model and forecast the El Niño phenomenon. The solar forcing is the only possible explanation for these results. Within their error limits, the detected three El Niño signals could represent the Hale cycle [24], the solar cycle[46, 56] and half the solar cycle. These strict El Niño periodicities indicate that the planets could cause a stationary and deterministic solar cycle [28]. A stochastic and non-stationary solar cycle [8, 54] could not cause the detected regular El Niño periodicities.
@@ Acknowledgements
Author contributions statement
L.J. peformed this research and wrote the manuscript.
@@
The author grants permission for publication.
@@
Main manuscript analysis: All data files and all DCM control files are stored to
https://zenodo.org/uploads/17018676.
Appendix analysis: All files, DCM Python code, DCM control files and Python code are stored to the
https://zenodo.org/records/18595333.
@@
Supplementary material
The $\mathcal{M}$= 1 analysis in Table S1 can be repeated using these two commands
cp longC14K110.dat dcm.dat
python dcm.py
The whole El Niño analysis can be repeated in similar fashion.
@@
Figure S1: First half forecast (Table S2: $\mathcal{M}$=2): Forecasting data (closed black circles) and forecasted data (open black circles). Green and red dottted lines denote one and three sigma model $g(t)$ errors. Residuals (closed and open blue circles) are offset to level of -3.
@@
Figure S2: First half forecast (Table S2: $\mathcal{M}$=3). Otherwise as in Figure S1.
@@
Figure S3: All data model $\mathcal{M}$=3 in Table S3. Otherwise as in Figure S1.
@@
Figure S4: All data model $\mathcal{M}$=4 in Table S3. Otherwise as in Figure S1.
Table S1: Trend in all data. Weighted DCM analysis performed between $P_{\mathrm{min}}=3$ and $P_{\mathrm{max}}=100$. (1) $\mathcal{M}$model, (2) $\eta$ free parameters and $\chi^{2}$ of model, (3) First signal: $P_{1}$ period, $A_{1}$ peak to peak amplitude and $t_{\mathrm{min},1}$ primary minimum epoch. (4-6) Next signals. (7-8) Fisher-test: $F_{\chi}$ test statistic and $Q_{\mathrm{F}}$ critical level. (9) Control file. Arrows point to best model in Fisher-test comparison.
@@
$\mathcal{M}$$_{4,1,2,\chi^{2}}$ 15 $0.34\pm 0.94~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{AD}}}$ $0.42\pm 0.90~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{AD}}}$ $0.50\pm 0.11$ $0.398\pm 0.084$ - - longC41K412.dat
@@
Table S2: Periods in first half data. Otherwise as in Table S1.
@@
Table S3: Periods in all weighted data. Otherwise as in Table S1.
@@
Table S4: Forecasts (Figures S3 and S4: open circles).
Figure S3 Figure S4
$t$ $y$ $\sigma$ $y$ $\sigma$
@@
Table S5: Periods in all non-weighted data. Otherwise as in Table S1.

### v6

@@ 3.11.3 Stability ( $\mathrm{C}_{3}$)
The artificial bootstrap data sets (Equation 31) represent “small changes in the input data”, while the bootstrap results for the model parameters represent “small chances in the solution” (Section 1: $\mathrm{C}_{3}$ condition formulation). We routinely check the stability of these solutions [30, Figure S5]. The unstable models, where the model parameter changes are large, are rejected (Section 2.1: $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}” models).
@@ 3.12 Forecast
There are numerous techniques for forecasting a time series [e.g., 23, 28, 33]. The DCM model $g(t_{i},\bm{\beta})$ can be used to forecast. We divide all data into the forecasting data and the forecasted data. The time points, the observations and the errors of these samples are
@@
This “Forecast-test” technique revealed at least five real signals in the sunspot record [30]
@@
during the chosen sample window $\Delta T^{\prime}=t_{n}^{\prime}-t_{1}^{\prime}$. Jetsu [30] used this $m_{\mathrm{Fore}}$ parameter to postdict the known $\Delta T^{\prime}$ time intervals of prolonged solar activity minima, like the Maunder minimum between the years 1640 and 1720. Since the known mean level of all sunspot data was $m$, Jetsu [30] used these three criteria for correct postdictions of past prolonged activity minima $\Delta T^{\prime}$ time intervals:
@@
The analyses of all seven complex time series indicate that the relative accuracy of amplitude and period estimates are lower than the relative accuracy of signal minimum and maximum epoch estimates (Tables 1 - 7). This statistical effect is the same when DCM is applied to the time series of any arbitrary phenomenon. This effect would, for example, explain why our solar cycle amplitude forecasts are less accurate than our solar cycle minimum and maximum epoch forecasts [30]. DCM detects the correct simulated period values, but DFT detects less accurate period values, which are not always correct. The leakage of DFT spectral power [32, 20] would explain why correct signal periods have not been detected earlier from the sunspot data.
@@ 4 Discussion
DCM proceeds through two stages. It computes the $R$ (Equation 14) or $\chi^{2}$ (Equation 15) values for a massive number of LS fits (Equation 50). The Gauß-Markov theorem ensures that the DCM model having the lowest $R$ or $\chi^{2}$ is inevitably always found, and the result for the non-linear iteration is unique (Equation 22). This first stage devours CPU. The second fast stage, the Fisher-test, then reveals the best DCM model alternative (Equation 48). Practical time series analysis applications, like digital signal processing, require fast computational algorithms [32]. Long computation time is the main AL constraining DCM. For example, Jetsu [30] computed about 4.5 million LS fits to search for four signals in the monthly sunspot data $(n=3287,K_{1}=4,K_{2}=1,K_{3}=0,n_{\mathrm{L}}=100,n_{\mathrm{S}}=30,n_{\mathrm{B}}=20)$. This required several months of CPU. The parallel Python code computations took a few days. Nevertheless, the CPU used is beside the point if a well-posed computational solution can be found to any challenging scientific ill-posed problem.
@@
A1 Appendix
Introduction
The modelling and forecasting of El Niño is considered notoriously difficult [55, 42, 6, 26, 35, 38, 53, 43]. We will show that our DCM can easily model and forecast the El Niño time series. Since the challenging analyses of all simulated extremely complex time series succeed (Sections 3.1-3.7), there are no rational reasons why our analysis of the simple El Niño time series would fail.
Data
Our El Niño data are from The Seasonal Mean Nino 4 HadISST1.1 (NOAA PSL) sample. These NOAA data were retrieved on January 29th, 2026. The twelve monthly sea surface temperature anomaly values give
  • $t_{i}=$ year
  • $y_{i}=$ mean of monthly values
  • $\sigma_{i}=$ standard deviation of monthly values
There were no values for the last four months of 2025. The full data contains $n=155$ yearly values between 1870 and 2024.
Results
The DCM can model and forecast the El Niño time series. This result may be the “holy grail” of climatology.
Trend
@@ Results
We detect significant signals only from the weighted data. Therefore, the trend and the forecasts are solved from the weighted data. The weighted DCM analysis results for all data are given in Table A1. For one, two, three, and four signals, the best trend alternative is the linear trend $K_{3}=1$. We use this $K_{3}=1$ trend in our DCM analysis of El Niño data. For the $\mathcal{M}$=3 model, the positive linear trend is $2M_{1}=0.56\pm 0.20$ $\mathrm{C}^{o}$ during $\Delta T=154$ years (Equation 5). This trend is approximately the same in all $\mathcal{M}$=1-4 models.
Forecasting data
@@
We use the first half of all data $(n=78)$ to forecast the second half $(n=77)$. The results of weighted DCM analysis are given in Table A2. The periods of the three strongest signals are $P_{1}=5.580\pm 0.085$, $P_{2}=12.82\pm 0.40$ and $P_{3}=19.30\pm 0.83$ years (Model $\mathcal{M}$=3). The forecasts of the two and three signals models, open circles, are shown in Figures A1 and A2. For the second half of the data, the forecast $\chi^{2}$ values for the one, two, three and four signal models are 421, 336, 370 and 428, respectively. Hence, the two signal $\mathcal{M}$=2 model gives the best forecast.
All data
@@
The DCM analysis results for all weighted data are given in Table A3. We detect the same periods that are already detected from the first half of all data. The three signal $\mathcal{M}$=3 model is the best. The El Niño signal periods are $P_{1}=5.662\pm 0.077$, $P_{2}=12.78\pm 0.12$ and $P_{3}=21.3\pm 1.5$ years. The DCM periodogram of the $\mathcal{M}$=3 model shows three clearly defined minima (Figure A3a). These three signals have the peak to peak amplitudes $A_{1}=0.40\pm 0.13$ $\mathrm{C}^{o}$, $A_{2}=0.47\pm 0.10$ $\mathrm{C}^{o}$ and $A_{3}=0.37\pm 0.12$ $\mathrm{C}^{o}$ (Figure A3b). The four signal $\mathcal{M}$=4 model is unstable $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}})$. The three and four signal models are shown in Figures A4 and A5. For both alternatives, the predicted values after 2024, open circles, are given in Table A4. Both model alternatives give essentially the same forecast.
@@
For the sake of consistency, we also give the results of non-weighted analysis of all data (Table A5). We detect the same signals from these non-weighted data, but the signal significances $Q_{\mathrm{F}}$ are lower. This confirms that the data error information of weighted data is crucial.
For all weighted data, the detection of the first signal is absolutely certain. The Fisher-test comparison between the $K_{3}=0$ pure noise model $g_{0,1,0,\chi^{2}}$ and the one signal model $g_{1,1,1,\chi^{2}}$ gives signal significance $Q_{\mathrm{F}}=3.9\times 10^{-13}$ (Table A6, $\mathcal{M}$=1). All other polynomial model alternatives, $1\leq K_{3}\leq 7$, are also rejected. The second and third signal significances are $Q_{\mathrm{F}}=2.7\times 10^{-7}$ and $Q_{\mathrm{F}}=0.00038$ (Table A3: $\mathcal{M}$=2-3).
Holy grail: El Niño, Sun and Planets
The DCM can model and forecast the El Niño phenomenon. The first half data forecast fits to the second half data (Figures A1b-d and A2b-d: open circles). For the two times larger all data sample, time will tell if our next decade forecast is also correct (Figures A4d and A5d: open circles). For example, we forecast an extreme El Niño between 2030 and 2032.
The solar forcing is the most probable explanation for these results. The three detected El Niño signal periods are close to the 22 years Hale cycle period [25], the 11 years solar cycle period [49, 59] and half the solar cycle period. A stochastic and non-stationary solar cycle [8, 57] can not cause these strict El Niño periodicities. Only very regular solar energy output changes can cause these El Niño signals because the noise added by the terrestrial climate response can not prevent our signal detection from the sea surface temperature data. These strictly periodic El Niño signals support the claim that the planets cause a stationary, multi-periodic and deterministic solar cycle(-s) [30].
Our mathematical model can detect regularities that the current physical models fail to detect, let alone forecast. Mathematics detects signals that should not be there according to Physics. The counterargument is that simple mathematics can not explain a complex physical phenomenon. We argue that the El Niño physical model is simple. The solar energy output determines El Niño. It is no surprise that the Sun warms the largest battery on the Earth, the Pacific Ocean. And if many cycles moderate the solar energy output, it is no surprise either that we have failed to discover the interference of these cycles, El Niño, in our yearly climate temperatures.
The current El Niño ”tool kit” is admittedly impressive [6]. Generations of hard work has given us those tools. Masses of scientists develop and use them. Yet, even the best currently available physical El Niño forecasts fail in 1.5 years [62]. Our mathematical DCM forecast can connect seven El Niño decades to the next seven El Niño decades (Figures A1 and A2). That is quite a coincidence if the chaos of physical models takes over in 1.5 years. The annual cost of El Niño damage is roughly one trillion dollars [9, 34]. An accurate El Niño prediction for just one year ahead would save the global economy trillions by mitigating long-term productivity losses [27, 61]. This should motivate other scientists to check our results immediately. [^4] All necessary file are in
https://zenodo.org/records/18726181/
No supercomputers are needed because an ordinary PC can detect the first two signals in about one minute. The third signal detection takes a few minutes.
William of Ockham (1287–1347) stated that if two competing hypotheses make the same predictions, the simplest hypothesis is usually the best one (Occam’s razor). The connection between El Niño, the Sun and the Planets may seem like the holy grail. However, those three are just physical phenomena. The real scientific holy grail is the DCM.
@@ Acknowledgements
Author contributions
L.J. did this research and wrote the manuscript.
@@
The author grants permission for publication in …
@@
All data files and Python codes are stored to
https://zenodo.org/records/18726181/
The manual for repeating all analysis is also published there.
@@
Figure A1: First half forecast (Table A2: $\mathcal{M}$=2): Forecasting data (closed black circles) and forecasted data (open black circles). Green and red dottted lines denote one and three sigma model $g(t)$ errors. Residuals (closed and open blue circles) are offset to level of -3. Units are x-axis: $t_{i}]$=y, y-axis: $y_{i}]=$ $\mathrm{C}^{o}$ and error bars = $1\times\sigma_{i}]=$ $\mathrm{C}^{o}$.
@@
Figure A2: First half forecast (Table A2: $\mathcal{M}$=3). Otherwise as in Figure A1.
Figure A3: All data model $\mathcal{M}$=3 in Table A3. (a) Periodogram (Equation 21). Otherwise as in Figures 2ab. (b) Three $h_{1}(t)$, $h_{2}(t)$ and $h_{3}(t)$ signals as a function of time $t$ (left-hand panels) and phase $\phi$ (right-hand panels) [30, Equation 29].
@@
Figure A4: All data model $\mathcal{M}$=3 in Table A3. Otherwise as in Figure A1.
@@
Figure A5: All data model $\mathcal{M}$=4 in Table A3. Otherwise as in Figure A1.
Table A1: Trend in all weighted data. DCM analysis between $P_{\mathrm{min}}=3$ and $P_{\mathrm{max}}=100$ years. Otherwise as in Table 8
@@
$\mathcal{M}$$_{4,1,2,\chi^{2}}$ 15 $0.34\pm 0.94~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{AD}}}$ $0.42\pm 0.90~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{AD}}}$ $0.50\pm 0.11$ $0.398\pm 0.084$ - - longC14K412.dat
@@
Table A2: Periods in first half data. Otherwise as in Table A1.
@@
Table A3: Periods in all weighted data. Otherwise as in Table A1.
@@
Table A4: Forecasts (Figures A4 and A5: open circles).
Figure A4 Figure A5
$t$ $y$ $\pm 1\sigma$ $y$ $\pm 1\sigma$
@@
Table A5: Periods in all non-weighted data. Otherwise as in Table A1.
@@
Table A6: Significance one signal model $\mathcal{M}$=1 in Table A3. Otherwise as in Table 9. 44footnotetext: Note: No test for $\mathcal{M}$=5  $(\eta_{1}=\eta_{2}=5)$.
(1) (2)
$g_{1,1,1}$
Polynomial $\eta=5,\chi^{2}=437$
$\mathcal{M}$=1 $\uparrow$
$g(t)=p(t,K_{3}=0)$ $F=19.9$
$\eta=1$, $\chi^{2}=670$ $Q_{\mathrm{F}}=3.9\times 10^{-13}$
$\mathcal{M}$=2 $\uparrow$
$g(t)=p(t,K_{3}=1)$ $11.4$
$\eta=2$, $\chi^{2}=537$ $Q_{\mathrm{F}}=9.2\times 10^{-7}$
$\mathcal{M}$=3 $\uparrow$
$g(t)=p(t,K_{3}=2)$ $F=17.0$
$\eta=3$, $\chi^{2}=537$ $Q_{\mathrm{F}}=2.1\times 10^{-7}$
$\mathcal{M}$=4 $\uparrow$
$g(t)=p(t,K_{3}=3)$ $F=26.5$
$\eta=4$, $\chi^{2}=515$ $Q_{\mathrm{F}}=7.8\times 10^{-7}$
$\mathcal{M}$=5 $\uparrow$
$g(t)=p(t,K_{3}=4)$ No test11footnotemark: 1
$\eta=5$, $\chi^{2}=514$ $Q_{\mathrm{F}}=1$
$\mathcal{M}$=6 $\uparrow$
$g(t)=p(t,K_{3}=5)$ F=22F=-22
$\eta=6$, $\chi^{2}=514$ $Q_{\mathrm{F}}=1$
$\mathcal{M}$=7 $\uparrow$
$g(t)=p(t,K_{3}=6)$ F=8.2F=-8.2
$\eta=7$, $\chi^{2}=492$ $Q_{\mathrm{F}}=1$
$\mathcal{M}$=8 $\uparrow$
$g(t)=p(t,K_{3}=7)$ F=5.4F=-5.4
$\eta=8$, $\chi^{2}=492$ $Q_{\mathrm{F}}=1$
[…]